The OHM RGB Floods have been screwed down in a fan pattern above the classroom door. Contrary to their wiki page, they're tested and working. Someone had removed the cable from one unit but it has been repaired. All channels working. PSU replaced, as Sun server one was loud enough to drown out all conversation in the classroom.
RPI and OLA installation tested and working with loaned Enttec USB Pro. Permanent installation stalled due to kernel module issues with cheap RS485 boards from ebay, now resolved.
Tested DMX PAR cans. One has been gutted and looks beyond repair, other working fine.
- Rebuild RPI using Ansible.
- Permanently install RPI above classroom doors.
- Investigate gutted DMX PAR properly.
- Install Working DMX PAR in classroom.
- Extend four core cable attached to each RGB flood. Position around classroom properly. Board uses Molex SL connectors.
- Install OLA either on babbage or babbages replacement. Port old ?beacons and ?dawn scripts to talk to OLA.
|01-03||RGB Flood 01|
|04-06||RGB Flood 02|
|07-09||RGB Flood 03|
|10-12||RGB Flood 04|
|13-15||RGB Flood 05|
|16-18||RGB Flood 06|
|19-21||RGB Flood 07|
|22-24||RGB Flood 08|
|25-27||RGB Flood 09|
|28-30||RGB Flood 10|